### v1

# Refuting Perfect Matchings in Spectral Expanders is Hard

Abstract
This work studies the complexity of refuting the existence of a perfect matching in spectral expanders with an odd number of vertices, in the Polynomial Calculus (PC) and Sum of Squares (SoS) proof system. Austrin and Risse [SODA, 2021] showed that refuting perfect matchings in sparse $d$-regular random graphs, in the above proof systems, with high probability requires proofs with degree $\Omega(n/\log n)$. We extend their result by showing the same lower bound holds for all $d$-regular graphs with a mild spectral gap.
1 Introduction
Perhaps the most fundamental problem in computation is to provide an answer to the question asked by Cook and Reckchow in their seminal paper [cook1979relative] – “Given a true statement $A$, is there a short proof of the claim that $A$ is true?” In trying to answer this question, we must first describe what constitutes a valid proof. That is, we must describe the language in which the proof is written (axioms), and the rules for checking it (the verifier). Each set of rules for writing and checking a proof defines a proof system. Therefore, a precise restatement of the question above is the following: “Given a true statement $A$ and a proof system $S$, what is the length of a shortest $\pi\in S$ that proves $A$?” If we could show that there exists a proof system $S$, such that for any true statement $A$, the length of the shortest proof in $S$ is upper bounded by some polynomial in the length of $A$, it would imply that $\mathsf{coNP}=\mathsf{NP}$ and consequently $\mathsf{P}=\mathsf{NP}$. Conversely, if we could show large proof size lower bounds for some $A$ in all proof systems, it would lead to a formal proof of the widely believed conjecture that $\mathsf{P}\neq\mathsf{NP}$. Unfortunately, such lower bounds for arbitrary proof systems are out of reach. As an intermediate step, the research community has invested a significant effort in proving lower bounds for increasingly expressive proof systems (e.g. see [abascal2021strongly, alekhnovich2001lower, atserias2020Size, buss1999linear, conneryd2023graph, de2023clique, impagliazzo1999lower, kothari2017SumOfSquares, potechin:LIPIcs.CCC.2020.38, raz2008elusive, razborov1998lower, schoenebeck2008linear]).
In this work, we focus on the algebraic and semi-algebraic proof systems of polynomial calculus (PC) and sum of squares (SoS). In algebraic proof systems, we are given a set $\mathcal{Q}=\{q_{i}(\mathbf{x})\text{ }|\text{ }i\in[m]\}$ of $m$ polynomial equations [^1] Semi-algebraic proof systems also allow for inequalities but we will not deal with inequality constraints in this paper. over $n$ variables $\mathbf{x}=\{x_{1},\dots,x_{n}\}$. In PC, the coefficientcs can be over an arbitrary field $\mathbb{F}$ and in the SoS the coefficients are over the reals. We say a proof $\pi$ is a refutation of $\mathcal{Q}$, if it is a proof of the claim (in the specified language) that there exists no assignment of $\mathbf{x}\in\mathbb{F}^{n}$ that satisfies all the polynomial equations in $\mathcal{Q}$. In PC and SoS, the proof $\pi$ is itself expressed as a sequence of polynomials. Two common measures of the complexity of the proof are the length of the sequence and the largest degree of the proof polynomials that refute $Q$ (see Section 2.1 for more formal details). Trade-offs between the two are well known; in particular, a lower bound on the degree implies a comparatively strong lower bound on the size of the proof. Therefore, in this work we focus on the former. We denote the smallest maximum degree over all proofs that refute $\mathcal{Q}$ in PC and SoS, with $\text{Deg}\left(Q\vdash_{\texttt{PC}_{F}}\bot\right)$ and $\text{Deg}\left(Q\vdash_{\texttt{SOS}}\bot\right)$, respectively. One motivation for proving lower bounds for algebraic proof systems, as opposed to propositional proof systems, is that often they imply lower bounds for a broad family of related algorithms for solving combinatorial optimisation problems. Similarly, upper bounding the proof length has led to the fruitful discovery of many efficient algorithms. The SoS proof system is of particular interest because of its close connection to the sum-of-squares hierarchy of semi-definite programming. We refer the reader to the survey by Fleming, Kothari, and Pitassi [fleming2019semialgebraic] for more details about the connections between the semi-algebraic proof systems and combinatorial optimisation.
In this work, we study the complexity of refuting perfect matchings in PC and SoS. Apart from being a natural problem in its own right, perfect matchings are also related to the pigeon hole principle [maciel2000new, pitassi1993exponential, raz2004resolution, razbarov2002pgp, razborov2003resolution] and Tseitin formula [filmus2013towards, galesi2019bounded, glinskih2017satisfiable, grigoriev2001linear], two well studied formulae in proof complexity. Assuming at most one pigeon fits in a single hole, the pigeon hole principle says $m$ pigeons cannot fit in $n<m$ holes. If we construct the complete bipartite graph with the left vertices as $m$ pigeons and the right vertices as $n<m$ holes, proving the pigeon hole principle amounts to proving that such bipartite graph does not have a perfect matching. There are other formulations of the pigeon hole principle (see the survey by Razborov [razbarov2002pgp]), and almost all of them have short proofs in the sum of squares proof system. In contrast, Tseitin formulae is known to require long proofs. The Tseitin formula over a graph claims that there is a spanning subgraph in which every vertex has odd degree. If a graph has a perfect matching, then the subgraph described by the matching ensures that every vertex has odd degree. However, formally refuting Tseitin formulae for expander graphs with odd number of vertices, in the SoS proof system, requires degree linear in the number of vertices in the graph [grigoriev2001linear]. Given its close connections to the pigeon-hole and Tseitin, and the different behaviour of the two formulae, it is natural to determine the complexity of refuting perfect matchings for non-bipartite graphs.
To refute perfect matchings in an algebraic proof system, we first need to specify combinatorial constraints as algebraic equalities. Given an undirected graph $G=(V,E)$, $V=\{1,\ldots,n\}$, and a vector $\mathbf{b}=(b_{1},\ldots,b_{n})\in\mathbb{F}^{n}$, we define $\text{Card}(G,\mathbf{b})$ as the following set of polynomial constraints over variables $x_{e}$ for $e\in E$:
[…]
For every $e\in E$, the equation $x_{e}(1-x_{e})=0$ restricts the domain of the above variables to bits. In plain words, $\text{Card}(G,\mathbf{b})$ denotes the claim that there exists a spanning subgraph $G^{\prime}\subseteq G$ such that a vertex $v\in V(G)$ has $b_{v}$ edges incident to it in $G^{\prime}$. Note if there was an assignment of variables $(x_{e})_{e\in E}$ that satisfies all equations in $\text{Card}(G,\mathbf{1})$, where $\mathbf{1}=(1,\ldots,1)\in\mathbb{F}^{n}$, it would imply that the graph $G$ has a perfect matching (given by the edges corresponding to variables with assignment 1). Therefore, we define $\text{PM}(G):=\text{Card}(G,\mathbf{1})$. When $|V|$ is odd, $G$ trivially does not contain a perfect matching. How difficult is it to refute $\text{PM}(G)$ in this case? In recent work, Austrin and Risse [Austrin_2022] showed that refuting $\text{PM}(G)$, in the Polynomial Calculus and Sum-of-Squares system, in the case $G$ is a random $d$-regular graphs with an odd number of vertices typically requires proofs with degree $\Omega\left(n/\log n\right)$. They conjecture (see [Austrin_2022, Section 6]) that the hardness results should also apply to general expander graphs but leave showing so as an open problem. In this work, we verify this by extending their result to all $d$-regular spectral expanders, that is, $d$-regular graphs with a mild condition on the spectral gap. In fact, similar to Austrin and Risse, we reduce the hardness of refuting $\text{Card}(G,\mathbf{t})$, for any odd value $t$, to the hardness of refuting $\text{Card}(G,\mathbf{1})$. As another special case, this answers the even-colouring case when $t=d/2$ is odd, a problem posed by Buss and Nordström [buss2021proof, Open Problem 7.7], which asks, “Are even colouring formulas over expander graphs hard for polynomial calculus over fields of characteristic distinct from 2 ?” Formally, we prove the following (for the definition of $(n,d,\lambda)$-graphs see Section 2.2).
Theorem 1.1 (Hardness Result For $\text{Card}(G,\mathbf{t})$)
There exist universal constants $\varepsilon,n_{0},d_{0}\in\mathbb{N}$ such that for any odd $n\geq n_{0}$ and even $d\in[d_{0},n]$, the following holds for any $(n,d,\lambda)$-graph $G$ with $\lambda<\varepsilon d$, and for any odd $1\leq t\leq d$:
$\text{Deg}\left(\text{Card}(G,\mathbf{t})\vdash_{\texttt{PC}_{F}}\bot\right)=\Omega\left(\nicefrac{{n}}{{\log n}}\right)$
$\text{Deg}\left(\text{Card}(G,\mathbf{t})\vdash_{\texttt{SOS}}\bot\right)=\Omega\left(\nicefrac{{n}}{{\log n}}\right)$
Proof overview.
We follow the overall approach of Austrin and Risse [Austrin_2022]. Very briefly, the strategy is to obtain an affine restriction (see Definition 2.6) $\text{Card}(G,\mathbf{t})|_{\rho}\equiv\text{PM}(H)$ where $H$ is some graph for which refuting $\text{PM}(H)$ requires large degree. An example of such $H$ is given by Buss, Grigoriev, and Impagliazzo [buss1999linear]. We now describe how to find such a restriction in more details: Using a result of Draganić, Krivelevich, and Nenadov [draganic22rolling], we show that $H$ topologically embeds into a given expander graph $G$ with $\lambda\leq\varepsilon d$ for some universal small constant $\varepsilon\in(0,1)$, such that all paths corresponding to the embedding have odd length. The main technical ingredient of Austrin and Risse is also similar embedding theorem, but one that critically relies on the underlying graph being random. As a random $d$-regular graph is with high probability an $(n,d,\lambda)$-graph with $\lambda=\Theta(\sqrt{d})$ (see [tikhomirov2016spectralgapdenserandom, Theorem A]), our embedding theorem subsumes that of [Austrin_2022]. Moreover, we show that one can find such an embedding so that the subgraph of $G$ induced by vertices which are not part of the embedding has a perfect matching. This allows us to use the restriction argument to transfer the hardness of $\text{PM}(H)$ into the hardness of $\text{PM}(G)$. To extend this to hardness of $\text{Card}(G,\mathbf{t})$ for an odd $3\leq t\leq d$, it suffices to show that the graph $G^{\prime}$ obtained from $G$ by removing all edges that participate in the embedding and the matching contains a $(t-1)$-regular spanning subgraph. Austrin and Risse achieve this using the contiguity property of random regular graphs. Instead, we provide a significantly simpler and shorter argument based on Tutte’s criterion.
The rest of the document as is structured as follows. In section 2, we describe the requisite background from graph theory and proof complexity required. In Section 3, we describe the machinery for finding a desired topological embedding. In Section 4, we prove conditions under which the residual graph has a perfect matching or, more generally, a $(t-1)$-regular spaning subgraph. In Section LABEL:sec:main-proof, we use the tools from the previous sections to prove Theorem 1.1. In Section LABEL:sec:related-work we briefly discuss a few other lower bounds using embeddings in proof complexity, and conclude with some future directions.
2 Preliminaries
2.1 Proof Complexity Preliminaries
Let $\mathcal{Q}=\{p_{1}=0,\dots,p_{m}=0\}$ be a set of polynomial equations [^2] The sum of squares proof system is a semi-algebraic proof system where $\mathcal{Q}$ may also contain inequalities of the form $p_{i}(x)\geq 0$. However, we only need equality constraints to express the existence of Perfect Matchings over graphs. Therefore to simplify our exposition, we write all our definitions using equality constraints only., which we refer to as axioms, over variables $\mathbf{X}=\{x_{1},\dots,x_{n},\bar{x}_{1},\dots,\bar{x}_{n}\}$.
Definition 2.1 (Sum Of Squares Refutation)
Given a set of $m$ polynomial equality constraints $\mathcal{Q}$ over the reals, a Sum of Squares (SoS) refutation is a set of polynomials $\pi=(t_{1},\dots,t_{m};s_{1},\dots,s_{a})$ such that
$h=\sum_{i\in[m]}t_{i}p_{i}+\sum_{i\in[a]}s_{i}^{2}=-1$
The degree of a proof $\pi$ is
$\text{Deg}\left(\pi\right)=\max\left\{\max_{i\in[m]}\text{Deg}\left(t_{i}\right)+\text{Deg}\left(p_{i}\right),\max_{i\in[a]}2\text{Deg}\left(s_{i}\right)\right\}$
Note that $s_{i}^{2}(x)\geq 0$ for any $x$ by definition. Therefore, if there were to exist some $x^{\star}$ such that $p_{i}(x^{\star})=0$ for all $p_{i}\in\mathcal{Q}$, then $\sum_{i\in[m]}t_{i}(x^{\star})p_{i}(x^{\star})=0$. This would imply that $h\geq 0$, but our proof $\pi$ shows that $h=-1$. Therefore, the existence of the set of polynomials $\pi$ act as a formal proof of the claim that the set of polynomials equations in $\mathcal{Q}$ is unsatisfiable.
Definition 2.2 (Complexity Of SoS Refutation)
If we let $\Pi$ denote the set of all valid SoS refutations for $\mathcal{Q}$, then the complexity of refuting $\mathcal{Q}$ in the SoS proof system is given by
$\text{Deg}\left(\mathcal{Q}\vdash_{\texttt{SOS}}\bot\right)=\min_{\pi\in\Pi}\text{Deg}\left(\pi\right)$
Polynomial Calculus (PC) is a dynamic version of the static Nullstellensatz proof system [fleming2019semialgebraic, Section 1.3] based on the following inference rules.
  1. 1.
    From polynomial equations $f=0$ and $g=0$ where $f,g\in\mathbb{F}[\mathbf{X}]$ we can derive $\alpha f+\beta g=0$ for $\alpha,\beta\in\mathbb{F}$.
  2. 2.
    From polynomial $f=0$ where $f\in\mathbb{F}[\mathbf{X}]$, we can derive $xf=0$ where $x\in\mathbf{X}$.
Definition 2.3 (Polynomial Calculus Refutations)
A Polynomial Calculus (PC) refutation of $\mathcal{Q}$ over is a set of polynomials $\pi=t_{1},\dots,t_{l}$ such that $t_{l}=1$, and for each $i\neq l$, either (1) $t_{i}\in\mathcal{Q}$, or (2) $t_{i}$ is derived from $\{t_{j}\}_{j<i}$ using the above rules. The degree of the proof is given by $\text{Deg}\left(\pi\right)=\max_{i\in l}\text{Deg}\left(t_{i}\right)$. If we let $\Pi$ denote the set of all PC refutations of $\mathcal{Q}$, then
$\text{Deg}\left(\mathcal{Q}\vdash_{\texttt{PC}_{F}}\bot\right)=\min_{\pi\in\Pi}\text{Deg}\left(\pi\right)$
To ensure boolean variables, we assume the axioms $\mathcal{Q}$ always contain the equations $x_{i}^{2}-x_{i}=0$ and $\bar{x}^{2}_{i}-\bar{x}_{i}=0$ for all $i\in[n]$. Equivalently, we can also just work in the ring $\mathbb{F}[x_{1},\dots,\bar{x}_{n}]/(x_{1}^{2}-x_{1},\dots,\bar{x}_{n}^{2}-\bar{x}_{n})$ of multi-linear polynomials. Multi-linearity implies that the degree of any proof can be at most $n$ i.e a proof of degree $\Omega\left(n\right)$ is the largest lower bound one can hope to achieve. Additionally, we will also assume that $01-x_{i}-\bar{x}_{i}=$ 0 is also included in $\mathcal{Q}$, for all $i\in[n]$, which ensures that the bar elements are bit complements of the non-bar elements. The following lemma is by Buss, Grigoriev, and Impagliazzo [buss1999linear] and gives an instance where perfect matching is hard to refute in the worst case.
Lemma 2.4 (Worst Case Hard Instance For PC)
Given any odd $n\in\mathbb{N}$, there exists a graph $H$ with $n$ vertices and maximum degree $\Delta_{H}=5$ such that Polynomial Calculus over any field of characteristic different from 2 requires degree $\Theta(n)$ to refute $\text{Card}(H,\mathbf{1})$.
A description of the worst case hard instance for SoS can be found in [Austrin_2022, Theorem A.3].
Lemma 2.5 (Worst Case Hard Instance For SOS)
Given any odd $n\in\mathbb{N}$, there exists a graph $H$ with $n$ vertices and maximum degree $\Delta_{H}=5$ such that SoS refutations requires degree $\Theta(n)$ to refute $\text{Card}(H,\mathbf{1})$.
2.1.1 Affine restriction
An important lemma we will need is that given a set of axioms $\mathcal{Q}$ over the ring $\mathbb{F}[x_{1},\dots,x_{n}]$, a partial assignment of variables can only make refuting $\mathcal{Q}$ easier. Given a set of $m$ polynomial equality constraints $\mathcal{Q}$ over boolean variables $\{x_{1},\dots,x_{n}\}$, let the family of functions $\{f_{i}:\{0,1\}^{n}\rightarrow\{\texttt{True},\texttt{False}\}\}_{i\in[m]}$, denote predicates for satisfiability for each constraint. For example, given $\alpha\in\{0,1\}^{n}$, $f_{i}(\alpha)=\texttt{True}$ if the $i$’th polynomial constraint $q_{i}\in\mathcal{Q}$ is satisfied i.e $p_{i}(\alpha)=0$. We say $\mathcal{Q}$ is satisfied if $\exists\alpha\in\{0,1\}^{n}$ such that $f_{i}(\alpha)=\texttt{True}\iff q_{i}(\alpha)=0$ for all $i\in[m]$. Given a map $\rho:\{x_{1},\dots,x_{n}\}\rightarrow\{x_{1},\dots,x_{n},\bar{x}_{1},\dots,\bar{x}_{n},1,0\}$, the restriction of a function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, denoted by $f|_{\rho}$, is defined as $f|_{\rho}(x_{1},\dots,x_{n})=f(\rho(x_{1}),\dots,\rho(x_{n}))$. Similarly, the restriction of formula $\mathcal{Q}$ is defined as $\mathcal{Q}|_{\rho}=\{f_{1}|_{\rho},\dots,f_{m}|_{\rho}\}$. Two formula $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are equivalent if they are element-wise equal, ignoring any functions that are constantly True. For example, $\mathcal{Q}=\{f_{a},f_{b},\texttt{True}\}$ and $\mathcal{Q}^{\prime}=\{f_{a},f_{b}\}$ are equivalent, denoted as $\mathcal{Q}\equiv\mathcal{Q}^{\prime}$.
Definition 2.6 (Affine Restriction)
We say that an axiom $\mathcal{Q}^{\prime}$ is an affine restriction of $\mathcal{Q}$ if there is a map $\rho:\{x_{1},\dots,x_{n}\}\rightarrow\{x_{1},\dots,x_{n},\bar{x}_{1},\dots,\bar{x}_{n},1,0\}$ such that $\mathcal{Q}\equiv\mathcal{Q}|_{\rho}$.
Lemma 2.7
Let $\mathcal{Q},\mathcal{Q}^{\prime}$ be axioms such that $\mathcal{Q}^{\prime}$ is an affine restriction of $\mathcal{Q}$, and each axiom of $\mathcal{Q}$ depends on a constant number of variables, then
  1. 1.
    For any field $\mathbb{F}$ it holds that $\text{Deg}\left(\mathcal{Q}\vdash_{\texttt{PC}_{F}}\bot\right)\in\Omega\left(\text{Deg}\left(\mathcal{Q}^{\prime}\vdash_{\texttt{PC}_{F}}\bot\right)\right)$
  2. 2.
    $\text{Deg}\left(\mathcal{Q}\vdash_{\texttt{SOS}}\bot\right)\in\Omega\left(\text{Deg}\left(\mathcal{Q}^{\prime}\vdash_{\texttt{SOS}}\bot\right)\right)$
The proof for the above lemma can be found in [Austrin_2022, Lemma 2.2]. What the above lemma says is that if we have a graph $G$ with odd vertices woth constant degree, that has a perfect matching on a subset of even vertices on the graph, then the size of the proof to refute $\text{PM}(G)$ is at least as large as refuting a perfect matching in $G$ with the even vertices removed.
2.2 Graph Theory Preliminaries
We use standard graph theoretic notation. For a graph $G$, we use $V(G)$ and $E(G)$ to denote the vertices and edges of $G$. For a vertex $v\in V(G)$, we use $\Gamma_{G}\left(v\right)=\{u\in V(G):(u,v)\in E^{\prime}\}$ to denote the neighbourhood of $v$ in $G$, and $\mathsf{deg}_{G}(v):=|\Gamma_{G}\left(v\right)|$. Given two sets $S,T\subseteq V(G)$, we use $e_{G}(S,T)$ to denote the number of edges in $G$ with one endpoint in $S$ and one endpoint in $T$. Note that we do not require $S$ and $T$ to be disjoint; in case they are not disjoint, every edge with both endpoints in $S\cap T$ is counted twice in $e_{G}(S,T)$. If the graph $G$ is clear from the context, we omit the subscript. Given $W\subseteq V(G)$, we denote with $G[W]$ the subgraph of $G$ induced by $W$. We say that a subgraph $G^{\prime}\subseteq G$ is spanning if $V(G^{\prime})=V(G)$.
Next, we give a definition of pseudorandom graphs.
Definition 2.8 ($(n,d,\lambda)$-graphs)
Let $G$ be a $d$-regular graph on $n$ vertices, and, let $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n}$ denote eigenvalues of the adjacency matrix of $G$. We say $G$ is an $(n,d,\lambda)$-graph if $\lambda(G):=\underset{{\{2,\dots,n\}}}{\max}|\lambda_{i}|\leq\lambda$.
The following is a well known result of Alon and Chung [alon88mixing].
Lemma 2.9 (Expander Mixing Lemma)
Given an $(n,d,\lambda)$-graph $G$, for any $S,T\subseteq V(G)$ we have
$\left|e_{G}(S,T)-\frac{d}{n}|S||T|\right|\leq\lambda\sqrt{|S||T|}$
We make use of the following two well known criteria of Tutte [tutte1952factors, tutte1947factorization]. Note that both of the lemmata ask for properties which are stronger than what Tutte criteria requires [^3] More specifically, in the Tutte criterion $q(G)$ denotes the number of odd sized connected components.; however, they are easier to state and verify in our application. We denote with $q(G)$ the number of connected components in a graph $G$.
Lemma 2.10 (Tutte’s Criterion)
If a graph $G$ has even number of vertices and for every subset $S\subseteq G$ we have $q(G\setminus S)\leq|S|$, then $G$ contains a perfect matching.
Lemma 2.11 (Tutte’s Generalised Criterion)
Let $f\in\mathbb{N}$ be even. Suppose $G$ is a graph such that for every pair of disjoint sets $S,T\subseteq V(G)$ the following holds:
$q(G\setminus(S\cup T))\leq|S|f-\sum_{w\in T}(f-|\Gamma_{G}\left(w\right)\setminus S|).$
Then $G$ contains a spanning subgraph $G^{\prime}\subseteq G$ which if $f$-regular.
2.3 Probabilistic Tools
Lemma 2.12 (Multiplicative Chernoff bound)
Suppose $X_{1},...,X_{n}$ are identical independent random variables taking values in $\{0,1\}$. Let $X$ denote their sum and let $\mu=n\mathbb{E}\left[X_{1}\right]$ denote the sum’s expected value. Then, for any $10<\delta<$ 1, we have
$\Pr\left[|X-\mu|\geq\delta\mu\right]\leq 2\exp(-\delta^{2}\mu/3)$
A dependency graph for a set of events $E_{1},...,E_{n}$ is a graph $G=(V,E)$ such that $V=\{1,...,n\}$ and, for $i=1,\dots,n$, event $E_{i}$ is mutually independent of the events $\{E_{j}|(i,j)\notin E\}$. The degree of the dependency graph is the maximum degree of any vertex in the graph.
Lemma 2.13 (Lovàsz Local Lemma)
Let $E_{1},...,E_{n}$ be a set of events, and assume that the following hold:
  1. 1.
    The degree of the dependency graph given by $(E_{1},\dots,E_{n})$ is bounded by $d$.
  2. 2.
    For all $i\in[n]$, $\Pr_{\mathcal{D}}\left[E_{i}\right]\leq\beta$
  3. 3.
    $\beta\leq\frac{1}{4d}$.
Then
[…]
The proof of (Multiplicative Chernoff bound) and (Lovàsz Local Lemma) can be found in any textbook on randomised algorithms (for example, see [mitzenmacher2017probability, Chapter 1, Chapter 7]). The following is the main result of this section.
Lemma 2.14
For every $10<c<$ 1 and $\gamma>0$, there exists $d_{0}$ such that the following holds. If $G$ is a $d$-regular graph, for some $d\geq d_{0}$, then there exists a subset $A\subseteq V(G)$ such that
$cd-\gamma d\leq\left|\Gamma_{G}\left(v\right)\cap A\right|\leq cd+\gamma d$
for every $v\in V(G)$.
Proof.
We prove the existence of such a partition $A\subseteq V(G)$ using the probabilistic method. For each $v\in V(G)$, we toss an independent coin $X_{i}$ with bias $c$. We include $v$ in $A$ if and only if $X_{i}=1$. Thus, $\mathbf{X}=(X_{1},\dots,X_{n})\in\{0,1\}^{n}$ is a random variable that describes how we choose $A$. For any $v\in V$, let $Y_{v}=|\Gamma_{G}\left(v\right)\cap A|$ denote the random variable that counts the number of neighbours of $v$ in $A$. Define $\delta=\gamma/c$, and for every $v\in V$ let $E_{v}=\mathbbm{1}\left\{|Y_{v}-dc|\geq\delta cd\right\}$ denote the bad event that $v$ has too many or too few neighbours in $A$. Observe that the dependency graph of events $\{E_{v}\}_{v\in V}$ has maximum degree at most $d^{2}$ (only vertices at most two hops away from $v$ affect how many of $v$’s neighbours are in $A$; there at most $d^{2}$ such vertices). As $G$ is $d$-regular, $\mathbb{E}_{\mathbf{X}}\left[Y_{v}\right]=cd$. By the (Multiplicative Chernoff bound), for any $v\in V$ we have $\Pr_{\mathbf{X}}\left[E_{v}\right]\leq 2\exp(-\delta^{2}cd/3)=:\beta$. For $d$ sufficiently large we have $\beta\leq 1/(4d^{2})$, and so $\beta d^{2}\leq 1/4$. All the conditions of (Lovàsz Local Lemma) are satisfied, from which we conclude that, with positive probability, none of the bad events happen. This implies the desired $A\subseteq V(G)$ exists. ∎
3 Topological embedding
In this section we describe the topological embedding result of Draganić, Krivelevich, and Nenadov [draganic22rolling]. We start with a necessary definition.
Definition 3.1 (Sub-divisions)
Given a graph $H$ and a function $\sigma:E(H)\rightarrow\mathbb{N}$, the $\sigma$-subdivision of $H$, denoted by ${H}^{\sigma}$, is the graph obtained by replacing each edge in $E(H)$ with a path of length $\sigma(e)$ joining the end points of $e$ such that all these paths are mutually vertex disjoint, except at the end points.
If a graph $G$ contains $H^{\sigma}$ for some $\sigma:E(H)\to\mathbb{N}$, then we say $G$ contains $H$ as a topological minor. In our application, it will be important that we can control the parity of $\sigma(e)$. The following result follows directly from [draganic22rolling, Theorem 1].
Theorem 3.2 (Embedding Theorem)
For every $D\in\mathbb{N}$ there exist $\alpha,\xi,C>0$, such that the following holds. Suppose $G$ is a graph with $n$ vertices and $m\geq Cn$ edges such that for every pair of disjoint subsets $S,T\subseteq V(G)$ of size $|S|,|T|\geq\xi n$, we have
$\bigl{|}e_{G}(S,T)-|S||T|p\bigr{|}\leq\xi|S||T|p$ (1)
where $p=m/\binom{n}{2}$. Then $G$ contains $H^{\sigma}$, where $H$ is any graph with maximum degree at most $D$, $H^{\sigma}$ has at most $\alpha n$ vertices, and $\sigma(e)\geq\log n$ for every $e\in E(H)$.
When $G$ is an $(n,d,\lambda)$ graph, we will we make use of Theorem 3.2 to show that $G$ satisfies the required properties, thereby contains $H$ as a topological minor. This gives us the following corollary.
Corollary 3.3
For every $D\in\mathbb{N}$ there exist $d_{0},n_{0}\in\mathbb{N},\varepsilon,\alpha\in(0,1)$, such that the following holds. Suppose $G$ is an $(n,d,\lambda)$-graph where $d\geq d_{0}$, and $\lambda<\varepsilon d$, and $n\geq n_{0}$. Let $B\subseteq V(G)$ be a subset of size $|B|\geq n/20$, and $H$ is any graph with maximum degree at most $D$ and at most $\alpha\frac{n}{\log n}$ vertices. Then the induced sub-graph $G[B]$ contains $H^{\sigma}$ such that $\sigma(e)$ is odd for every $e\in E(H)$.
Proof.
Let $m$ denote the number of edges in the induced subgraph $G[B]$, which gives us $2m=e_{G}(B,B)$. Denote $b=|B|$ and Define $p=m/\binom{b}{2}$. By the (Expander Mixing Lemma), we have
$\left|2m-\frac{d}{n}b^{2}\right|\leq\lambda b.$
Dividing both sides with $b(b-1)$, and observing that $\frac{db^{2}}{n(b-1)b}=\frac{d}{n}\frac{b}{b-1}=\frac{d}{n}(1+\frac{1}{b-1})$, we further get
$\left|\frac{2m}{b(b-1)}-\frac{d}{n}-\frac{d}{n(b-1)}\right|$ $\leq\frac{\lambda}{b-1}$ (2)
From this we have
$\left|p-\frac{d}{n}\right|$ $=\left|p-\frac{d}{n}-\frac{d}{n(b-1)}+\frac{d}{n(b-1)}\right|$ (3)
$\leq\left|p-\frac{d}{n}-\frac{d}{n(b-1)}\right|+\frac{1}{(b-1)}$ (4)
$\leq\frac{\lambda}{(b-1)}+\frac{1}{(b-1)}$ (5)
$\leq\frac{2\lambda}{b}.$ (6)
Let us briefly justify each step: (4) comes from the triangle inquality and $d/n\leq 1$; Equation (5) comes from Equation (2); the last inequality comes from the fact that $b\in\Omega\left(n\right)$, so the inequality holds for $n$ large enough.
Let $\xi$ be as given by (Embedding Theorem). Using the bound on the difference between $p$ and $d/n$, for every disjoint subsets $S,T\subseteq B$ of size $|S|,|T|\geq\xi n$, for $\lambda<\varepsilon d$ where $\varepsilon$ is sufficiently small, we have
$\Bigl{|}e_{G}(S,T)-p|S||T|\Bigr{|}$ $\leq\left|e_{G}(S,T)-\frac{d}{n}|S||T|\right|+\left|\frac{d}{n}|S||T|-p|S||T|\right|$
$\leq\lambda\sqrt{|S||T|}+\frac{2\lambda}{b}|S||T|\leq\xi|S||T|p.$
With the lower bounds on $S,T$ and $B$, we can make $\varepsilon$ sufficiently small with respect to $\xi$ to get the upper bound in the last step. Let $\sigma:E(H)\to\mathbb{N}$ be the constant function where $\sigma(e)$ is the smallest odd integer larger than $\log n$. As $G[B]$ has at least $Cn$ edges (by (Expander Mixing Lemma)), $\sigma(e)\leq 2+\log n$, and $H$ has at most $\alpha\frac{n}{\log n}$ vertices, we can invoke (Embedding Theorem) to conclude that $G[B]$ contains $H^{\sigma}$. ∎
4 Perfect matching and regular subgraphs
As described earlier, the second ingredient in our hardness proof is showing that a certain residual graph contains a perfect matching or a spanning $(t-1)$-regular subgraph. In this section we state and prove these ingredients.
Lemma 4.1 (Perfect Matching Lemma)
Suppose $G$ is an $(n,d,\lambda)$-graph, where $\lambda<d/50$. Suppose $U\subseteq V(G)$ is such that $\delta(G[U])\geq\frac{9d}{10}$. Then for all $S\subseteq U$, $G[U\setminus S]$ has at most $|S|$ connected components, that is, $q(G[U\setminus S])\leq|S|$.
Proof.
Consider some $U\subseteq V(G)$ such that $\delta(G[U])\geq\frac{9d}{10}$. Consider some $S\subseteq U$. We aim to show that $G[U\setminus S]$ has at most $|S|$ connected components. If $|S|\geq|U|/2$ then as $G[U\setminus S]$ has fewer than $|S|$ vertices, so the the upper bound on connected components trivially holds. Thus, we can assume $|S|<|U|/2$. We claim the following:
If $\text{If }e_{G}(X,Y)\geq 1\text{ for every partition $X\cup Y=U\setminus S$, $|X|,|Y|\geq|S|/3$, then }q(G[U\setminus S])\leq|S|.$ (7)
To see why, assume towards a contradiction that there exists an edge between every partition $X\cup Y=U\setminus S$, where $|X|,|Y|\geq|S|/3$, and $G[U\setminus S]$ has more than $|S|$ connected components. Denote the vertex sets of these components by $C_{1},\ldots,C_{k}$, for some $k>|S|$. Let $X^{\star}:=C_{1}\cup\ldots\cup C_{s}$ and $Y^{\star}:=C_{s+1}\cup\ldots\cup C_{k}$, where $s=\lfloor|S|/2\rfloor\geq|S|/3$. By construction, even if each component $C_{i}$ is a singleton set, we get that $|X^{\star}|,|Y^{\star}|\geq|S|/3$. Now as all $C_{i}$’s are disjoint connected components, there can be no edge between $X^{\star}$ and $Y^{\star}$. Therefore, we have found a correct sized partition of $U\setminus S$ without an edge between them, which contradicts our assumption.